OpenStudy (anonymous):

How many gallons of bleach are in x gallons of a 70% bleach solution? 0.07x 0.7x 70x

OpenStudy (anonymous):

.7x is the answer 70% is 70 parts per 100 70/100 = 7/10 = .7

OpenStudy (anonymous):

thank you robz8

OpenStudy (anonymous):

yw